write an equation of the line containing the given point and perpendicular to the given line. Express your answer in the form y=mx+b. (8,9); 6x+y=2
Answer by Cromlix(4381)

6x+y=2
Put into y = mx +c form
y = -6x + 2
Gradient (m) =-6
If a line is perpendicular to another
their gradients multiply together to
make -1 (m1 x m2 = -1)
Therefore m2 = 1/6
Using formula y - b = m(x - a)
and coordinates (8,9)
y - 9 = 1/6(x - 8)
y - 9 =1/6x - 8/6
y = 1/6x - 8/6 + 9
y = 1/6x + 46/6
OR 6y = x + 46
Hope this helps.
:-)